OpenStudy (anonymous):

Damien wants to deposit $1200 into savings accounts at three different banks: Bank of Chance, Merchant Bank, and Utopian Financing. He will deposit three times as much into Merchant Bank as Bank of Chance because they offer a higher interest rate. He also expects the Utopian Financing deposit to be only 20% of the total of the other two deposits. How much will Damien deposit into the Utopian Financing savings account?

OpenStudy (anonymous):

Let's start by defining some variables: Let c = Amount deposited in Bank of Chance Let m = Amount deposited in Merchant Bank Let u = Amount deposited in Utopian Financing

OpenStudy (y2o2):

x+y+z = 1200 x = 3y z = 0.2(x+y) = 0.2(4x)

OpenStudy (anonymous):

Because we know the total amount is $1200, our first equation will be: c+m+u = 1200

OpenStudy (anonymous):

"He will deposit three times as much into Merchant Bank as Bank of Chance" How would that be interpreted?

OpenStudy (anonymous):

What you want to note here is that three times as much goes into Merchant Bank. So Merchant Bank will be 3 times what Bank of Chance will be. You can express that like this: 3c=m

OpenStudy (anonymous):

Finally: "He also expects the Utopian Financing deposit to be only 20% of the total of the other two deposits."

OpenStudy (anonymous):

So that means: u=20%(c+m)

OpenStudy (anonymous):

Now, we don't work with percents, so we need to turn that into a decimal. u=.2(c+m)

OpenStudy (anonymous):

So the equations we're left with are: c+m+u=1200 3c=m u=.2(c+m)

OpenStudy (anonymous):

From here, let's plug in what we know from the second and third equations into the first equation: c+m+u=1200 Well, we know that m=3c and we know that u=.2(c+m)

OpenStudy (anonymous):

So we're really looking at: c+(3c)+.2(c+(3c))=1200

OpenStudy (anonymous):

c+(3c)+.2(c+(3c))=1200 4c + .2(4c) = 1200 4c + .8c = 1200 4.8c=1200 /4.8 /4.8 c=250

OpenStudy (anonymous):

And, if we know c = 250 then we can figure out m, because m = 3c m=3(250) m=750

OpenStudy (anonymous):

And now that we know those two amounts and we know the total has to be 1200, we can use that first equation: c+m+u=1200 250+750+u=1200 1000+u=1200 -1000 -1000 u=200

OpenStudy (anonymous):

So finally: c=$250 m=$750 u=$200 And your teacher will probably want you to answer the question in a complete sentence. :)